The length of the rectangle is three times the width. The area is 48. Find the length, width, and perimeter
Answer by JBarnum(2146)

l=length w=width p=perimeter a=area
given:
lw=a
l=3w
a=48
p=2(l+w)
p=8w
.
substitution work:
48=3w(w)
48=3w^2
16=w^2
4=w
l=12
.
p=8(4)
p=32